# Determination of the density in the linear elastic wave equation

### v2

@@ 4. S-waves and proof of the main theorem
$+2(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}\frac{\lambda+3\mu}{\rho}c_{S}^{-1}g^{\beta\gamma}\partial_{\beta}c_{S}^{2}A_{S}\eta_{\gamma}-2\partial_{\alpha}\log\sqrt{\rho}c_{S}g^{\delta\beta}\partial_{\beta}c^{2}_{S}A_{S}\eta_{\delta}$
$-c_{S}\partial_{\gamma}c_{S}g^{\beta\gamma}\partial_{\beta}A_{S}c_{S}\eta_{\alpha}+c_{S}^{2}g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}\partial_{\beta}A_{S}c_{S}\eta_{\alpha}$
$+c_{S}^{-3}\partial_{3}c_{S}\mu c_{S}\partial_{3}A_{S}\eta_{\alpha}-\frac{1}{2}g^{\beta\gamma}c_{S}^{-2}\frac{\partial g_{\beta\gamma}}{\partial\tau}\mu c_{S}\partial_{3}A_{S}\eta_{\alpha}$
@@
$+2\partial_{3}\log\sqrt{\rho}g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}c_{S}A_{S}\eta_{\beta}+4c_{S}^{2}g^{\delta\gamma}\partial_{\delta}\log\sqrt{\rho}\Gamma^{\beta}_{\alpha\gamma}c_{S}A_{S}\eta_{\beta}$
$+4c_{S}g^{\beta\gamma}\partial_{\beta}c_{S}\partial_{\alpha}c_{S}A_{S}\eta_{\gamma}$
@@
$\begin{split}g^{\beta\gamma}\partial_{\beta}\partial_{\gamma}A_{S}=&J^{-1/2}c_{S}^{-1/2}g^{\beta\gamma}\partial_{\beta}\partial_{\gamma}\rho^{-1/2}+2g^{\beta\gamma}\partial_{\beta}J^{-1/2}c_{S}^{-1/2}\partial_{\gamma}\rho^{-1/2}+2J^{-1/2}g^{\beta\gamma}\partial_{\beta}c_{S}^{-1/2}\partial_{\gamma}\rho^{-1/2}\\ =&-J^{-1/2}c_{S}^{-1/2}\rho^{-1/2}g^{\beta\gamma}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}+J^{-1/2}c_{S}^{-1/2}\rho^{-1/2}g^{\beta\gamma}(\partial_{\alpha}\log\sqrt{\rho})(\partial_{\beta}\log\sqrt{\rho})\\ &+g^{\beta\gamma}J^{-3/2}\partial_{\beta}Jc_{P}^{-1/2}\rho^{-1/2}\partial_{\gamma}\log\sqrt{\rho}+J^{-1/2}\rho^{-1/2}g^{\beta\gamma}c_{S}^{-3/2}\partial_{\beta}c_{S}\partial_{\gamma}\log\sqrt{\rho}\\ &+\frac{1}{\sqrt{\rho}}\times\text{terms that do not depend on $\rho$},\end{split}$
@@
$H^{\prime}=$ $\partial^{2}_{3}\log\sqrt{\rho}\eta_{\alpha}-(\partial_{3}\log\sqrt{\rho})^{2}\eta_{\alpha}-c_{S}^{-1}\partial_{3}c_{S}\partial_{3}\log\sqrt{\rho}\eta_{\alpha}-{J^{-1}\partial_{3}J\partial_{3}\log\sqrt{\rho}\eta_{\alpha}}+c_{S}^{-2}\partial_{3}c_{S}^{2}\partial_{3}\log\sqrt{\rho}\eta_{\alpha}$
@@
$-2\partial_{3}\log\sqrt{\rho}\left(-\partial_{3}\log\sqrt{\rho}-\frac{1}{2}c_{S}^{-1}\partial_{3}c_{S}-{\frac{1}{2}J^{-1}\partial_{3}J}\right)\eta_{\alpha}+g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}\partial_{3}\log\sqrt{\rho}\eta_{\beta}$
@@
$+g^{\beta\gamma}\partial_{\beta}c_{S}^{2}\partial_{\gamma}\log\sqrt{\rho}\eta_{\alpha}-2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\left(-\partial_{\gamma}\log\sqrt{\rho}-\frac{1}{2}c_{S}^{-1}\partial_{\gamma}c_{S}-{\frac{1}{2}J^{-1}\partial_{\gamma}J}\right)\eta_{\alpha}$
$+c_{S}^{2}g^{\beta\gamma}\partial_{\gamma}\partial_{\beta}\log\sqrt{\rho}\eta_{\alpha}-c_{S}^{2}g^{\beta\gamma}(\partial_{\gamma}\log\sqrt{\rho})(\partial_{\beta}\log\sqrt{\rho})\eta_{\alpha}-{c_{S}^{2}g^{\beta\gamma}J^{-1}\partial_{\beta}J\partial_{\gamma}\log\sqrt{\rho}\eta_{\alpha}}$
$-c_{S}^{2}g^{\beta\gamma}c_{S}^{-1}\partial_{\beta}c_{S}\partial_{\gamma}\log\sqrt{\rho}\eta_{\alpha}$
$+g^{\beta\gamma}\partial_{\gamma}c_{S}^{2}\partial_{\alpha}\log\sqrt{\rho}\eta_{\beta}+2g^{\beta\gamma}c_{S}^{2}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}$
$+2(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}g^{\gamma\beta}\partial_{\beta}c_{S}^{2}\eta_{\gamma}+4(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}g^{\gamma\beta}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}$
$+\frac{2c_{P}^{2}+c_{S}^{2}}{c_{S}^{2}}\partial_{\alpha}(c_{S}^{2})g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}+\partial_{\alpha}(c_{P}^{2}-2c_{S}^{2})g^{\delta\varepsilon}\partial_{\varepsilon}\log\sqrt{\rho}\eta_{\delta}$
$-c_{S}^{2}\partial_{\alpha}g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}+2\partial_{\alpha}g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}+g^{\delta\gamma}c_{S}^{2}\Gamma^{\beta}_{\alpha\gamma}\partial_{\delta}\log\sqrt{\rho}\eta_{\beta}-g^{\beta\gamma}\Gamma^{\delta}_{\alpha\gamma}c_{S}^{2}\partial_{\delta}\log\sqrt{\rho}\eta_{\beta}$
$+(-2\partial_{\alpha}c_{S}^{2}-\frac{8c_{S}^{2}\partial_{\alpha}c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}})g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}$
$+2(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}(c_{P}^{2}+c_{S}^{2})c_{S}^{-1}g^{\beta\gamma}\partial_{\beta}c_{S}\eta_{\gamma}-2\partial_{\alpha}\log\sqrt{\rho}c_{S}g^{\delta\beta}\partial_{\beta}c_{S}\eta_{\delta}$
$+c_{S}\partial_{\gamma}c_{S}g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}\eta_{\alpha}-c_{S}^{2}g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}\partial_{\beta}\log\sqrt{\rho}\eta_{\alpha}$
$-c_{S}^{-3}\partial_{3}c_{S}c_{S}^{2}\partial_{3}\log\sqrt{\rho}\eta_{\alpha}+\frac{1}{2}g^{\beta\gamma}c_{S}^{-2}\frac{\partial g_{\beta\gamma}}{\partial\tau}c_{S}^{2}\partial_{3}\log\sqrt{\rho}\eta_{\alpha}-4c_{S}^{2}g^{\beta\gamma}\partial_{\beta}c_{S}\partial_{\alpha}\log\sqrt{\rho}\eta_{\gamma}$
+terms that do not depend on $+\text{terms that do not depend on $\rho$}.$
@@
$\begin{split}&\partial_{\alpha}g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}+g^{\delta\gamma}c_{S}^{2}\Gamma^{\beta}_{\alpha\gamma}\partial_{\delta}\log\sqrt{\rho}\eta_{\beta}-g^{\beta\gamma}\Gamma^{\delta}_{\alpha\gamma}c_{S}^{2}\partial_{\delta}\log\sqrt{\rho}\eta_{\beta}\\ =&c_{S}^{2}\left(-g^{\beta\delta}\partial_{\alpha}g_{\delta\varepsilon}g^{\varepsilon\gamma}+\frac{1}{2}g^{\beta\delta}g^{\gamma\varepsilon}(\partial_{\alpha}g_{\delta\varepsilon}+\partial_{\delta}g_{\alpha\varepsilon}-\partial_{\varepsilon}g_{\alpha\delta})-\frac{1}{2}g^{\gamma\delta}g^{\beta\varepsilon}(\partial_{\alpha}g_{\delta\varepsilon}+\partial_{\delta}g_{\alpha\varepsilon}-\partial_{\varepsilon}g_{\alpha\delta})\right)\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}\\ =&-c_{S}^{2}g^{\gamma\delta}g^{\beta\varepsilon}(\partial_{\alpha}g_{\delta\varepsilon}+\partial_{\delta}g_{\alpha\varepsilon}-\partial_{\varepsilon}g_{\alpha\delta})\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}\\ =&-2c_{S}^{2}g^{\gamma\delta}\Gamma^{\beta}_{\alpha\delta}\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}.\end{split}$
@@
$c_{S}^{-2}|\eta|^{2}=1,$
that is $g_{\alpha\beta}\eta^{\alpha}\eta^{\beta}=g^{\alpha\beta}\eta_{\alpha}\eta_{\beta}=c_{S}^{2}$. Using also (11) and (12), we then have
$c_{S}^{-2}H^{\prime}_{\alpha}\eta^{\alpha}$
$=$ $\partial^{2}_{3}\log\sqrt{\rho}+(\partial_{3}\log\sqrt{\rho})^{2}+c_{S}^{2}g^{\beta\gamma}(\partial_{\gamma}\log\sqrt{\rho})(\partial_{\beta}\log\sqrt{\rho})+c_{S}^{2}g^{\beta\gamma}\partial_{\gamma}\partial_{\beta}\log\sqrt{\rho}+$
@@
$+2\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}\eta^{\beta}\eta^{\alpha}+4(c_{P}^{2}-2c_{S}^{2})\frac{1}{c_{P}^{2}-c_{S}^{2}}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}\log\sqrt{\rho}\eta^{\beta}\eta^{\alpha}-c_{S}^{2}g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}\partial_{\beta}\log\sqrt{\rho}$
@@
$+\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}\partial_{3}\log\sqrt{\rho}+c_{S}^{-2}\frac{\partial g_{\alpha\gamma}}{\partial\tau}\partial_{3}\log\sqrt{\rho}\eta^{\gamma}\eta^{\alpha}-2\Gamma^{\beta}_{\alpha\delta}\partial_{\beta}\log\sqrt{\rho}\eta^{\delta}\eta^{\alpha}$
$+c_{S}^{-1}\partial_{3}c_{S}\partial_{3}\log\sqrt{\rho}+3c_{S}g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}\partial_{\gamma}c_{S}$
$+3\frac{c_{P}^{2}-c_{S}^{2}}{c_{S}^{4}}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}+c_{S}^{-2}\partial_{\alpha}c_{P}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}-\frac{12}{c_{P}^{2}-c_{S}^{2}}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}$
+terms that do not depend on $+\text{terms that do not depend on $\rho$}$
$=$ $\partial^{2}_{3}\log\sqrt{\rho}-c_{S}^{-1}\partial_{3}c_{S}\partial_{3}\log\sqrt{\rho}+c_{S}g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}\partial_{\gamma}c_{S}+c_{S}^{2}g^{\beta\gamma}\partial_{\gamma}\partial_{\beta}\log\sqrt{\rho}$
$+\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}\partial_{3}\log\sqrt{\rho}-c_{S}^{2}g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}\partial_{\beta}\log\sqrt{\rho}$
$+2\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}\eta^{\beta}\eta^{\alpha}+c_{S}^{-2}\frac{\partial g_{\alpha\gamma}}{\partial\tau}\partial_{3}\log\sqrt{\rho}\eta^{\gamma}\eta^{\alpha}-2\Gamma^{\beta}_{\alpha\delta}\partial_{\beta}\log\sqrt{\rho}\eta^{\delta}\eta^{\alpha}$
$+c_{S}^{2}g^{\beta\gamma}(\partial_{\gamma}\log\sqrt{\rho})(\partial_{\beta}\log\sqrt{\rho})+(\partial_{3}\log\sqrt{\rho})^{2}$
$+4(c_{P}^{2}-2c_{S}^{2})\frac{1}{c_{P}^{2}-c_{S}^{2}}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}\log\sqrt{\rho}\eta^{\beta}\eta^{\alpha}$
$+c_{S}^{-2}\partial_{3}c_{S}^{2}\partial_{3}\log\sqrt{\rho}+g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}\partial_{\gamma}c_{S}^{2}$
$+3\frac{c_{P}^{2}-c_{S}^{2}}{c_{S}^{4}}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}+c_{S}^{-2}\partial_{\alpha}c_{P}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}-\frac{12}{c_{P}^{2}-c_{S}^{2}}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}$
+terms that do not depend on $+\text{terms that do not depend on $\rho$}$
$=$ $c_{S}^{2}\Delta\log\sqrt{\rho}+2\nabla_{\alpha}\nabla_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}+c_{S}^{2}|\nabla\log\sqrt{\rho}|^{2}+4\frac{c_{P}^{2}-2c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}\nabla_{\alpha}\log\sqrt{\rho}\nabla_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}$

### v3

@@ Appendix A Proof of lemmas
$+2(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}\frac{\lambda+3\mu}{\rho}c_{S}^{-1}g^{\beta\varepsilon}\partial_{\beta}c_{S}^{2}A_{S}\eta_{\varepsilon}-2\partial_{\alpha}\log\sqrt{\rho}c_{S}g^{\delta\beta}\partial_{\beta}c^{2}_{S}A_{S}\eta_{\delta}$
$-c_{S}\partial_{\varepsilon}c_{S}g^{\beta\varepsilon}\partial_{\beta}A_{S}c_{S}\eta_{\alpha}+c_{S}^{2}g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}\partial_{\beta}A_{S}c_{S}\eta_{\alpha}$
$+c_{S}^{-3}\partial_{3}c_{S}\mu c_{S}\partial_{3}A_{S}\eta_{\alpha}-\frac{1}{2}g^{\beta\varepsilon}c_{S}^{-2}\frac{\partial g_{\beta\varepsilon}}{\partial\tau}\mu c_{S}\partial_{3}A_{S}\eta_{\alpha}$
@@
$+2\partial_{3}\log\sqrt{\rho}g^{\beta\varepsilon}\frac{\partial g_{\alpha\varepsilon}}{\partial\tau}c_{S}A_{S}\eta_{\beta}+4c_{S}^{2}g^{\delta\varepsilon}\partial_{\delta}\log\sqrt{\rho}\Gamma^{\beta}_{\alpha\varepsilon}c_{S}A_{S}\eta_{\beta}$
$+4c_{S}g^{\beta\varepsilon}\partial_{\beta}c_{S}\partial_{\alpha}c_{S}A_{S}\eta_{\varepsilon}$
@@
$\begin{split}g^{\beta\varepsilon}\partial_{\beta}\partial_{\varepsilon}A_{S}=&J^{-1/2}c_{S}^{-1/2}g^{\beta\varepsilon}\partial_{\beta}\partial_{\varepsilon}\rho^{-1/2}+2g^{\beta\varepsilon}\partial_{\beta}J^{-1/2}c_{S}^{-1/2}\partial_{\varepsilon}\rho^{-1/2}+2J^{-1/2}g^{\beta\varepsilon}\partial_{\beta}c_{S}^{-1/2}\partial_{\varepsilon}\rho^{-1/2}\\ =&-J^{-1/2}c_{S}^{-1/2}\rho^{-1/2}g^{\beta\varepsilon}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}+J^{-1/2}c_{S}^{-1/2}\rho^{-1/2}g^{\beta\varepsilon}(\partial_{\alpha}\log\sqrt{\rho})(\partial_{\beta}\log\sqrt{\rho})\\ &+g^{\beta\varepsilon}J^{-3/2}\partial_{\beta}Jc_{P}^{-1/2}\rho^{-1/2}\partial_{\varepsilon}\log\sqrt{\rho}+J^{-1/2}\rho^{-1/2}g^{\beta\varepsilon}c_{S}^{-3/2}\partial_{\beta}c_{S}\partial_{\varepsilon}\log\sqrt{\rho}\\ &+\frac{1}{\sqrt{\rho}}\times\text{terms that do not depend on $\rho$},\end{split}$
@@
$H^{\prime}_{\alpha}=$ $\partial^{2}_{3}\log\sqrt{\rho}\eta_{\alpha}-(\partial_{3}\log\sqrt{\rho})^{2}\eta_{\alpha}-c_{S}^{-1}\partial_{3}c_{S}\partial_{3}\log\sqrt{\rho}\eta_{\alpha}-{J^{-1}\partial_{3}J\partial_{3}\log\sqrt{\rho}\eta_{\alpha}}+c_{S}^{-2}\partial_{3}c_{S}^{2}\partial_{3}\log\sqrt{\rho}\eta_{\alpha}$
@@
$-2\partial_{3}\log\sqrt{\rho}\left(-\partial_{3}\log\sqrt{\rho}-\frac{1}{2}c_{S}^{-1}\partial_{3}c_{S}-{\frac{1}{2}J^{-1}\partial_{3}J}\right)\eta_{\alpha}+g^{\beta\varepsilon}\frac{\partial g_{\alpha\varepsilon}}{\partial\tau}\partial_{3}\log\sqrt{\rho}\eta_{\beta}$
@@
$+g^{\beta\varepsilon}\partial_{\beta}c_{S}^{2}\partial_{\varepsilon}\log\sqrt{\rho}\eta_{\alpha}-2g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\left(-\partial_{\varepsilon}\log\sqrt{\rho}-\frac{1}{2}c_{S}^{-1}\partial_{\varepsilon}c_{S}-{\frac{1}{2}J^{-1}\partial_{\varepsilon}J}\right)\eta_{\alpha}$
$+c_{S}^{2}g^{\beta\varepsilon}\partial_{\varepsilon}\partial_{\beta}\log\sqrt{\rho}\eta_{\alpha}-c_{S}^{2}g^{\beta\varepsilon}(\partial_{\varepsilon}\log\sqrt{\rho})(\partial_{\beta}\log\sqrt{\rho})\eta_{\alpha}-{c_{S}^{2}g^{\beta\varepsilon}J^{-1}\partial_{\beta}J\partial_{\varepsilon}\log\sqrt{\rho}\eta_{\alpha}}$
$-c_{S}^{2}g^{\beta\varepsilon}c_{S}^{-1}\partial_{\beta}c_{S}\partial_{\varepsilon}\log\sqrt{\rho}\eta_{\alpha}$
$+g^{\beta\varepsilon}\partial_{\varepsilon}c_{S}^{2}\partial_{\alpha}\log\sqrt{\rho}\eta_{\beta}+2g^{\beta\varepsilon}c_{S}^{2}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}\eta_{\varepsilon}$
$+2(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}g^{\varepsilon\beta}\partial_{\beta}c_{S}^{2}\eta_{\varepsilon}+4(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}g^{\varepsilon\beta}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta_{\varepsilon}$
$+\frac{2c_{P}^{2}+c_{S}^{2}}{c_{S}^{2}}\partial_{\alpha}(c_{S}^{2})g^{\beta\varepsilon}\partial_{\beta}\log\sqrt{\rho}\eta_{\varepsilon}+\partial_{\alpha}(c_{P}^{2}-2c_{S}^{2})g^{\delta\varepsilon}\partial_{\varepsilon}\log\sqrt{\rho}\eta_{\delta}$
$-c_{S}^{2}\partial_{\alpha}g^{\beta\varepsilon}\partial_{\beta}\log\sqrt{\rho}\eta_{\varepsilon}+2\partial_{\alpha}g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta_{\varepsilon}+g^{\delta\varepsilon}c_{S}^{2}\Gamma^{\beta}_{\alpha\varepsilon}\partial_{\delta}\log\sqrt{\rho}\eta_{\beta}-g^{\beta\varepsilon}\Gamma^{\delta}_{\alpha\varepsilon}c_{S}^{2}\partial_{\delta}\log\sqrt{\rho}\eta_{\beta}$
$+(-2\partial_{\alpha}c_{S}^{2}-\frac{8c_{S}^{2}\partial_{\alpha}c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}})g^{\beta\varepsilon}\partial_{\beta}\log\sqrt{\rho}\eta_{\varepsilon}$
$+2(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}(c_{P}^{2}+c_{S}^{2})c_{S}^{-1}g^{\beta\varepsilon}\partial_{\beta}c_{S}\eta_{\varepsilon}-2\partial_{\alpha}\log\sqrt{\rho}c_{S}g^{\delta\beta}\partial_{\beta}c_{S}\eta_{\delta}$
$+c_{S}\partial_{\varepsilon}c_{S}g^{\beta\varepsilon}\partial_{\beta}\log\sqrt{\rho}\eta_{\alpha}-c_{S}^{2}g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}\partial_{\beta}\log\sqrt{\rho}\eta_{\alpha}$
$-c_{S}^{-3}\partial_{3}c_{S}c_{S}^{2}\partial_{3}\log\sqrt{\rho}\eta_{\alpha}+\frac{1}{2}g^{\beta\varepsilon}c_{S}^{-2}\frac{\partial g_{\beta\varepsilon}}{\partial\tau}c_{S}^{2}\partial_{3}\log\sqrt{\rho}\eta_{\alpha}-4c_{S}^{2}g^{\beta\varepsilon}\partial_{\beta}c_{S}\partial_{\alpha}\log\sqrt{\rho}\eta_{\varepsilon}$
$+\text{terms that do not depend on $\rho$}.$
@@
$\begin{split}&\partial_{\alpha}g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta_{\varepsilon}+g^{\delta\varepsilon}c_{S}^{2}\Gamma^{\beta}_{\alpha\varepsilon}\partial_{\delta}\log\sqrt{\rho}\eta_{\beta}-g^{\beta\varepsilon}\Gamma^{\delta}_{\alpha\varepsilon}c_{S}^{2}\partial_{\delta}\log\sqrt{\rho}\eta_{\beta}\\ =&c_{S}^{2}\left(-g^{\beta\delta}\partial_{\alpha}g_{\delta\varepsilon}g^{\varepsilon\kappa}+\frac{1}{2}g^{\beta\delta}g^{\kappa\varepsilon}(\partial_{\alpha}g_{\delta\varepsilon}+\partial_{\delta}g_{\alpha\varepsilon}-\partial_{\varepsilon}g_{\alpha\delta})-\frac{1}{2}g^{\kappa\delta}g^{\beta\varepsilon}(\partial_{\alpha}g_{\delta\varepsilon}+\partial_{\delta}g_{\alpha\varepsilon}-\partial_{\varepsilon}g_{\alpha\delta})\right)\partial_{\beta}\log\sqrt{\rho}\eta_{\kappa}\\ =&-c_{S}^{2}g^{\kappa\delta}g^{\beta\varepsilon}(\partial_{\alpha}g_{\delta\varepsilon}+\partial_{\delta}g_{\alpha\varepsilon}-\partial_{\varepsilon}g_{\alpha\delta})\partial_{\beta}\log\sqrt{\rho}\eta_{\kappa}\\ =&-2c_{S}^{2}g^{\varepsilon\delta}\Gamma^{\beta}_{\alpha\delta}\partial_{\beta}\log\sqrt{\rho}\eta_{\varepsilon}.\end{split}$
@@
$c_{S}^{2}|\eta|^{2}=c^{2}_{S}g^{\alpha\beta}\eta_{\alpha}\eta_{\beta}=c^{2}_{S}g_{\alpha\beta}\eta^{\alpha}\eta^{\beta}=1.$
Using also (16) and (17), we then have
$\langle H^{\prime},\eta\rangle_{g_{S}}=c_{S}^{2}g^{\alpha\beta}H^{\prime}_{\alpha}\eta_{\beta}$
$=$ $\partial^{2}_{3}\log\sqrt{\rho}+(\partial_{3}\log\sqrt{\rho})^{2}+c_{S}^{2}g^{\beta\varepsilon}(\partial_{\varepsilon}\log\sqrt{\rho})(\partial_{\beta}\log\sqrt{\rho})+c_{S}^{2}g^{\beta\varepsilon}\partial_{\varepsilon}\partial_{\beta}\log\sqrt{\rho}+$
$+2c_{S}^{4}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}+4(c_{P}^{2}-2c_{S}^{2})\frac{c_{2}^{4}}{c_{P}^{2}-c_{S}^{2}}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}\log\sqrt{\rho}\eta^{\beta}\eta^{\alpha}-c_{S}^{2}g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}\partial_{\beta}\log\sqrt{\rho}$
@@
$+\frac{1}{2}g^{\beta\varepsilon}\frac{\partial g_{\beta\varepsilon}}{\partial\tau}\partial_{3}\log\sqrt{\rho}+c_{S}^{2}\frac{\partial g_{\alpha\varepsilon}}{\partial\tau}\partial_{3}\log\sqrt{\rho}\eta^{\varepsilon}\eta^{\alpha}-2c_{S}^{4}\Gamma^{\beta}_{\alpha\delta}\partial_{\beta}\log\sqrt{\rho}\eta^{\delta}\eta^{\alpha}$
$+c_{S}^{-1}\partial_{3}c_{S}\partial_{3}\log\sqrt{\rho}+3c_{S}g^{\beta\varepsilon}\partial_{\beta}\log\sqrt{\rho}\partial_{\varepsilon}c_{S}$
$+3(c_{P}^{2}-c_{S}^{2})\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}+c_{S}^{2}\partial_{\alpha}c_{P}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}-\frac{12c_{S}^{4}}{c_{P}^{2}-c_{S}^{2}}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}$
$+\text{terms that do not depend on $\rho$}$
$=$ $\partial^{2}_{3}\log\sqrt{\rho}-c_{S}^{-1}\partial_{3}c_{S}\partial_{3}\log\sqrt{\rho}+c_{S}g^{\beta\varepsilon}\partial_{\beta}\log\sqrt{\rho}\partial_{\varepsilon}c_{S}+c_{S}^{2}g^{\beta\varepsilon}\partial_{\varepsilon}\partial_{\beta}\log\sqrt{\rho}$
$+\frac{1}{2}g^{\beta\varepsilon}\frac{\partial g_{\beta\varepsilon}}{\partial\tau}\partial_{3}\log\sqrt{\rho}-c_{S}^{2}g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}\partial_{\beta}\log\sqrt{\rho}$
@@
$+2c_{S}^{4}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}\eta^{\beta}\eta^{\alpha}+c_{S}^{2}\frac{\partial g_{\alpha\varepsilon}}{\partial\tau}\partial_{3}\log\sqrt{\rho}\eta^{\varepsilon}\eta^{\alpha}-c_{S}^{4}2\Gamma^{\beta}_{\alpha\delta}\partial_{\beta}\log\sqrt{\rho}\eta^{\delta}\eta^{\alpha}$
$+c_{S}^{2}g^{\beta\varepsilon}(\partial_{\varepsilon}\log\sqrt{\rho})(\partial_{\beta}\log\sqrt{\rho})+(\partial_{3}\log\sqrt{\rho})^{2}$
$+4(c_{P}^{2}-2c_{S}^{2})\frac{c_{S}^{4}}{c_{P}^{2}-c_{S}^{2}}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}\log\sqrt{\rho}\eta^{\beta}\eta^{\alpha}$
$+c_{S}^{2}\partial_{3}c_{S}^{2}\partial_{3}\log\sqrt{\rho}+g^{\beta\varepsilon}\partial_{\beta}\log\sqrt{\rho}\partial_{\varepsilon}c_{S}^{2}$
$+3(c_{P}^{2}-c_{S}^{2})\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}+c_{S}^{2}\partial_{\alpha}c_{P}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}-\frac{12c_{S}^{4}}{c_{P}^{2}-c_{S}^{2}}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}$
$+\text{terms that do not depend on $\rho$}$
$=$ $c_{S}^{2}\Delta\log\sqrt{\rho}+2c_{S}^{4}\nabla_{\alpha}\nabla_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}+c_{S}^{2}|\nabla\log\sqrt{\rho}|^{2}+4c_{S}^{4}\frac{c_{P}^{2}-2c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}\nabla_{\alpha}\log\sqrt{\rho}\nabla_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}$